The sum of the present ages of Eric and his father is 72 years. In 12 years, his father will be twice as old as Eric will be at that time. Find their present ages.
Answer by mananth(16946)

The sum of the present ages of Eric and his father is 72 years. In 12 years, his father will be twice as old as Eric will be at that time. Find their present ages.
Eric's age =x
fathers age = y
x+y =72...........................(1)
After 12 years
y+12= 2(x+12)
y+12=2x+24
y-2x=12...........................(2)


x+y =72...........................(1)
y-2x=12...........................(2)
multiply equation (1) by -1
we get -x-y=-72
Add this equation to (2)
-3x=-60
/3
x= 20
Eric's age = 20 years
Father's age = 72-20 = 52 years